# Transformers for dynamical systems learn transfer operators in-context

### v1

@@ Appendix B Model architecture and training
A Architecture
We implement a compact causal language model operating on a discrete vocabulary ($V=100$) that predicts the next token autoregressively. Our model is a decoder-only transformer, similar to GPT models but with relative positional embedding akin to the T5 family [7, 28]. Each input token is mapped to a higher-dimensional embedding ($d_{\text{model}}=256$). The network itself consists of two identical Transformer-style residual blocks, each using single-head causal self-attention, with query and key projections into $d_{k}=128$ dimensions, and a value projection back into $d_{\text{model}}$ [36]. Attention logits are scaled by $d_{k}^{-1/2}$ and strictly masked with a lower-triangular causal mask. Each block applies pre-normalization with LayerNorm before attention [4], then adds the attention output residually, and then applies a second LayerNorm followed by a position-wise feed-forward network with expansion factor $2$ (resulting in hidden width $2d_{\text{model}}=512$), using ReLU activation and a second linear layer returning to $d_{\text{model}}$, again added residually. A final LayerNorm precedes the output projection to logits over the vocabulary. The output projection is weight-tied to the input embedding matrix. The maximum context length is a fixed block size of $T_{\max}=512$ tokens. Positional information is encoded using ALiBi, an additive attention-logit bias proportional to the nonnegative lag distance with slope $\alpha=1.0$, i.e. a term $\alpha\,(i-j)$ for $i\geq j$ and 0 otherwise.
B Tokenization
We convert real-valued time series into tokens with mean scaling followed by uniform scalar quantization into $V$ bins, following the approach used in the first generation of the Chronos family of time series foundation models [3]. We favor direct quantization over patch-based approaches used in other time series foundation models due to the interpretability of the token space, and because it allows us to design our study based on prior studies of in-context generalization in language models [15].
@@ B Tokenization
and we then normalize the entire series $\tilde{x}_{i}=x_{i}/s$. The normalized values are discretized into integer bin IDs in $\{1,\dots,V\}$ using uniformly spaced bin centers $c_{j}$ spanning $c_{\min},c_{\max}]$ (with $c_{\min}$ and $c_{\max}$ treated as hyperparameters), and decision boundaries placed at midpoints between adjacent centers, with outer edges extending to $\pm\infty$ so that all values map to a unique bin. A reserved “padding” token uses ID 0. Missing values (NaNs) are handled by replacing their contribution with 0 for purposes of scale computation, while their token positions are set to the padding ID 0 in the emitted token sequence. Decoding maps bin IDs back to their corresponding centers and rescales via $\hat{x}_{i}=s\,\tilde{x}_{i}$.
C Training
We train our model using next-token prediction with cross-entropy loss over the vocabulary. We quantize the training time series and provide it to the model as a single long one-dimensional token stream, from which we sample fixed-length windows of length $T=512$ uniformly at random: the input is a contiguous block of $512$ tokens and the target is the same block shifted by one position. The effective dataset length is treated as a large virtual size ($10,000$ samples) with random window sampling each time. We optimize the model using AdamW with learning rate $5\times 10^{-5}$, weight decay $0.0$, batch size $32$, with gradient accumulation over micro-batches (with default size $1$) by averaging loss across micro-steps before backpropagation [22]. We clip gradients to a maximum global norm of $1.0$ at each step. We train for $50,000$ optimization steps and perform validation every $1000$ steps by evaluating cross-entropy on one randomly sampled validation batch, and additionally on a separate out-of-distribution validation stream of the same form. We do not use a learning-rate schedule.
D Inference
We perform autoregressive inference for a specified number of new tokens $H$ beyond the provided context. At each step, we use only the most recent $C=512$ context tokens. The model outputs logits for the next token, which are converted to a token choice either greedily (argmax, temperature 0) or by sampling. We append the sampled token to the context, and repeat the process for $H$ steps, yielding a lengthened token sequence. If a real-valued time series is needed for downstream analysis, we decode the output token sequence back to real values via dequantization to bin centers followed by multiplication by the stored mean scale $s$.
@@ Appendix C Estimating the fully-observed transfer operator with Ulam’s method
Partition and symbolic dynamics. We partition $\{B_{1},\dots,B_{K}\}$ the data $\{\mathbf{y}_{t}\}$ by clustering. Given a set of samples $\{\mathbf{y}_{t}\}$ we use $K$-means clustering to generate Voronoi cells $B_{i}\equiv\{\mathbf{y}:\arg\min_{j}\|\mathbf{y}-\mathbf{c}_{j}\|=i\}$ with centers $\{\mathbf{c}_{i}\}_{i=1}^{K}$, $\mathbf{c}_{i}\in\mathbb{R}^{D}$. This defines a (hard) coding map $\kappa:\mathbf{y}\to\{1,\dots,K\}$, which we apply to the training data to generate the symbolic sequence
@@
Empirical Ulam matrix at lag $\tau$. For a fixed lag $\tau$, Ulam’s method approximates the transfer operator $\mathcal{P}_{\tau}$ by the $K\times K$ matrix $P^{(\tau)}$ as
@@
i.e., the probability mass transferred from bin $B_{i}$ to bin $B_{j}$ over time $\tau$. We compute this quantity by counting all transition pairs, and dividing by the total number of transitions.
@@
Choice of partition size. Because Ulam discretization trades bias (too few bins) against variance/sparsity (too many bins), we scan over candidate $K$ and select a value that yields informative, non-degenerate transition structure. We compute a row-wise entropy statistic of $P^{(\tau)}$ across $K$ and choose values that maximize this statistic. This heuristic favors partitions that are neither so large that most transitions are self-loops, nor so small that the transition matrix becomes overly sparse.
Appendix D Estimating the manifold dimension of the Lorenz-96 system
@@
with cyclic boundary conditions $X_{i+n}=X_{i}$ and $X_{i-n}=X_{i}$. Here $X_{i}(t)$ are the state variables, $D$ is the system dimension, and $F$ is the constant external forcing. In all simulations we set $F=8.0$. The initial condition is taken as $X_{i}(0)=F$ for all $i$, with a random normal perturbation of $10^{-2}$ added to the first component to ensure diversity across generated trajectories. The system is integrated to $1000$ time units using the implicit Radau IIA method. We discard the first 500 steps of the generated trajectory as transients.
@@
Appendix E Attention rollout

### v2

@@ Appendix B Model architecture and training
Architecture.We implement a compact causal language model operating on a discrete vocabulary ($V=100$) that predicts the next token autoregressively. Our model is a decoder-only transformer, similar to GPT models but with relative positional embedding [5, 12]. Each input token is mapped to a higher-dimensional embedding ($d_{\text{model}}=256$). The network consists of two identical transformer-style residual blocks, each using single-head causal self-attention, with query and key projections into $d_{k}=128$ dimensions, and a value projection back into $d_{\text{model}}$ [13]. Attention logits are scaled by $d_{k}^{-1/2}$ and strictly masked with a lower-triangular causal mask. Each block applies pre-normalization with LayerNorm before attention [3], then adds the attention output residually, and then applies a second LayerNorm followed by a position-wise feed-forward network with expansion factor $2$ (resulting in hidden width $2d_{\text{model}}=512$), using ReLU activation and a second linear layer returning to $d_{\text{model}}$, again added residually. A final LayerNorm precedes the output projection to logits over the vocabulary. The output projection is weight-tied to the input embedding matrix. The maximum context length is a fixed block size of $C=512$ tokens. Positional information is encoded using ALiBi, an additive attention-logit bias equal to the nonnegative lag distance: (ij)-(i-j) for $i\geq j$ and 0 otherwise.
Tokenization. We convert real-valued time series into tokens with mean scaling followed by uniform scalar quantization into $V$ bins, following the approach used in the first generation of the Chronos family of time series foundation models [2]. We favor direct quantization over patch-based approaches used in other time series foundation models due to the interpretability of the token space, and because it allows us to design our study based on prior studies of in-context generalization in language models [8, 4].
@@
and we then normalize the entire series $\tilde{x}_{i}=x_{i}/s$. The normalized values are discretized into integer bin IDs in $\{1,\dots,V\}$ using uniformly spaced bin centers $c_{j}$ spanning $c_{\min},c_{\max}]$ (with $c_{\min}$ and $c_{\max}$ treated as hyperparameters), and decision boundaries placed at midpoints between adjacent centers, with outer edges extending to $\pm\infty$ so that all values map to a unique bin. A reserved “padding” token uses ID 0. Missing values (NaNs) are handled by replacing their contribution with 0 for purposes of scale computation, while their token positions are set to the padding ID 0 in the emitted token sequence.
Training. We train our model using next-token prediction with cross-entropy loss over the vocabulary. We quantize the training time series and provide it to the model as a single long one-dimensional token stream, from which we sample fixed-length windows of length $T=512$ uniformly at random: the input is a contiguous block of $512$ tokens and the target is the same block shifted by one position. The effective dataset length is treated as a large virtual size ($10,000$ samples) with random window sampling each time. We optimize the model using AdamW with learning rate $5\times 10^{-5}$, weight decay $0.0$, batch size $32$, with gradient accumulation over micro-batches (with default size $1$) [10]. We clip gradients to a maximum global norm of $1.0$ at each step. We train for $50,000$ optimization steps and perform validation every $1000$ steps by evaluating cross-entropy validation batches. We do not use a learning-rate schedule.
Inference. We perform autoregressive inference for a specified number of new tokens $H$ beyond the provided context. At each step, we use only the most recent $C=512$ context tokens. The model outputs logits for the next token, which are converted to a token choice via sampling. We append the sampled token to the context, and repeat the process for $H$ steps, yielding a lengthened token sequence.
@@ Appendix C Estimating the fully-observed transfer operator with Ulam’s method
Partitioning. We partition $\{B_{1},\dots,B_{K}\}$ the data $\{\mathbf{y}_{t}\}$ by clustering. Given a set of samples $\{\mathbf{y}_{t}\}$ we use $K$-means clustering to generate Voronoi cells $B_{i}\equiv\{\mathbf{y}:\arg\min_{j}\|\mathbf{y}-\mathbf{c}_{j}\|=i\}$ with centers $\{\mathbf{c}_{i}\}_{i=1}^{K}$, $\mathbf{c}_{i}\in\mathbb{R}^{D}$. This defines a coding map $\kappa:\mathbf{y}\to\{1,\dots,K\}$, which we apply to the training data to generate the symbolic dynamics
@@
Estimating the Ulam matrix. For a fixed lag $\tau$, Ulam’s method approximates the transfer operator $\mathcal{P}_{\tau}$ by the $K\times K$ matrix $P^{(\tau)}$ as
@@
i.e., the probability mass transferred from bin $B_{i}$ to bin $B_{j}$ over time $\tau$. We compute this quantity by counting all transition pairs separated by $\tau$, and dividing by the total number of transitions observed exiting the symbol associated with each row.
@@
Choosing the partition size. Because Ulam discretization trades bias (too few bins) against variance/sparsity (too many bins), we scan over candidate $K$ and select a value that yields informative, non-degenerate transition structure. We compute a row-wise entropy statistic of $P^{(\tau)}$ across $K$ and choose values that maximize this statistic. This heuristic favors partitions that are neither so large that most transitions are self-loops, nor so small that the transition matrix becomes overly sparse.
Appendix D Variation of model properties with context length
To probe the dependence of the model’s properties on context length, we retrain the model with varying context lengths $C$. We find that the epochwise double-descent phenomenon that we report in the main text becomes more pronounced at longer context lengths (Fig. S1), consistent with existing results for increasing model complexity on a finite training dataset [11]. Comparing the final training and validation losses across models, the training and in-distribution test losses both decrease with increasing context, though the latter plateaus. The out-of-distribution test loss plateaus early, implying that the simplistic transfer operator strategy employed by the model in-context does not benefit from additional context during pretraining. We expect that this occurs because only low-order $k$-gram statistics are needed to estimate the transfer operator.
Figure S1: (A) Test-OOD loss curves for replicate models trained with varying context lengths $C$. (B) The final cross-entropy loss for training data, in-distribution test data, and out-of-distribution test data, as a function of context length $C$.
Appendix E Estimating the manifold dimension of the Lorenz-96 system
@@
with cyclic boundary conditions $X_{i+D}=X_{i}$ and $X_{i-D}=X_{i}$. Here $X_{i}(t)$ are the state variables, $D$ is the system dimension, and $F$ is the constant external forcing. In all simulations we set $F=8.0$. The initial condition is taken as $X_{i}(0)=F$ for all $i$, with a random normal perturbation of $10^{-2}$ added to the first component to ensure diversity across generated trajectories. The system is integrated to $1000$ time units using the implicit Radau IIA method. We discard the first 500 steps of the generated trajectory as transients.
@@
Appendix F Attention rollout
@@
Appendix G In-context learning of transfer operators
Our results follow from a recent study that shows the ability of two-layer transformers to estimate Markov chains in-context [7]: in a dynamical systems setting, transfer operator estimation corresponds to bigram counting on delay coordinates. For a Markov-chain prediction task with a Dirichlet prior over transition rows, the Bayes-optimal in-context predictor is the Dirichlet-multinomial posterior mean. Prior work shows that a 2-layer single-head, attention-only transformer can implement the bigram-counting circuit needed to estimate this mean [7].
Let $x_{t}\in\{1,2,$...,\} be the quantized scalar time series. For an embedding length $m$, we define the delay state $s_{t}=(x_{t},x_{t-1},\dots,x_{t-m+1})\in\mathcal{S}$. We assume $m$ is large enough that the delay-state process is an approximately first-order Markov chain
$\Pr(s_{t+1}\mid s_{t},s_{t-1},\dots)\approx\Pr(s_{t+1}\mid s_{t})=P_{s_{t},s_{t+1}},$
where $P$ is a row-stochastic transition matrix on $\mathcal{S}$, which consists of $M$ distinct states. In order to model the dynamics, the transformer seeks to construct an approximation of this operator $\hat{P}$ given finite-length context.
For a given context, a one-step forecast reduces to estimating the row $p_{i}$ associated with the current delay state $s_{t}=i$.
$\Pr(s_{t+1}\mid s_{t}=i)=p_{i}.$
The model seeks to estimate this next-state conditional by counting transitions between pairs of delay states.
Assume row $p_{i}$ has a Dirichlet prior $p_{i}\sim\mathrm{Dir}(\alpha)$, and let $c_{ij}$ be the number of observed transitions $i\to j$ seen within the context window,
$N_{i}=\sum_{j}c_{ij}.$
Then the posterior mean is
$\widehat{P}_{ij}=\frac{c_{ij}+\alpha_{j}}{N_{i}+\sum_{j}\alpha_{j}}.$ (A4)
For large $N_{i}$, the Dirichlet prior contributes only minor corrections, so $\widehat{p}_{i}$ is asymptotically the empirical row-frequency estimator.
Scaling of KL divergence. Let the distribution $\mathbf{p}_{i}\equiv\{p_{ij}\}_{j=1}^{M}$ describe the $i^{th}$ row of the true transition matrix. The row $\mathbf{p}_{i}$ has an effective support size $M_{i}\leq M$, corresponding to the number of outgoing states reachable from state $i$. Let $\hat{\mathbf{p}}_{i}\equiv\{\hat{p}_{ij}\}_{j=1}^{M}$ correspond to an estimator of the true $\mathbf{p}_{i}$ derived from observing $N_{i}$ transitions in the context originating from state $s_{i}$. Assuming $\left|\hat{p}_{ij}-p_{ij}\right|\ll 1$ for all $j$, a second-order Taylor expansion of the KL divergence gives
$D_{\text{KL}}(\mathbf{p}_{i}\|\hat{\mathbf{p}}_{i})\approx\frac{1}{2}\sum_{j}\frac{(\hat{p}_{ij}-p_{ij})^{2}}{p_{ij}}.$ (A5)
The expected value of the KL divergence over many samplings of set $\hat{\mathbf{p}}_{i}$ is
$\mathbb{E}\!\left[D_{\text{KL}}(\mathbf{p}_{i}\|\hat{\mathbf{p}}_{i})\right]\approx\frac{1}{2}\sum_{j}\frac{\mathbb{E}\!\left[(\hat{p}_{ij}-p_{ij})^{2}\right]}{p_{ij}}.$ (A6)
We treat $\hat{\mathbf{p}}_{i}$ as a random variable with multinomial statistics due to count-based estimation. The variance of a given element is given by
[…]
We substitute this into Eq. A8
$\mathbb{E}\!\left[D_{\text{KL}}(\mathbf{p}_{i}\|\hat{\mathbf{p}}_{i})\right]\approx\frac{1}{2}\sum_{j}\frac{1-p_{ij}}{N_{i}}.$ (A8)
$\mathbb{E}\!\left[D_{\text{KL}}(\mathbf{p}_{i}\|\hat{\mathbf{p}}_{i})\right]\approx\frac{M_{i}-1}{2N_{i}}.$ (A9)
Thus the expected error of a single row of the estimated transition operator scales as $\mathcal{O}(N_{i}^{-1})$, where $N_{i}$ is the number of transitions originating from state $i$ observed in the context.
We assume the process $p_{ij}$ is stationary with invariant distribution $\bm{\pi}^{(0)}$, and mixing is strong enough that over a context of length $C$ the number of expected counts of each state is given by the invariant distribution,
$N_{i}\approx C\pi^{(0)}_{i}$.
Thus, once the delay embedding is large enough to make the process approximately Markov, increasing context length improves estimation of the transfer operator as $1/C$.
Scaling laws for the spectrum of the transfer operator. We describe approximated transfer operator $\hat{p}\in\mathbb{R}^{M\times M}$ as the ground truth transfer operator $p$ and an unknown error matrix $e\in\mathbb{R}^{M\times M}$, $\hat{p}=p+e$. We estimate rows of $\hat{p}$ based on $\mathcal{O}(C)$ effective samples overall, and so concentration gives
$\|E\|=\mathcal{O}(C^{-1/2}),$ (A10)
up to dimension and occupancy factors.
If $\lambda$ is an isolated eigenvalue of $P$, then Weyl’s bounds from eigenvalue perturbation theory correspond to
$|\hat{\lambda}-\lambda|=O(C^{-1/2}).$ (A11)
Likewise, if $\bm{\pi}^{(k)}\in\mathbb{R}^{D}$ denotes an eigenvector associated with the eigenvalue $\lambda_{i}$ separated by spectral gap $\mathrm{gap}_{i}=\min_{j\neq i}\left|\lambda_{i}-\lambda_{j}\right|$, then Davis-Kahan bounds imply
$\|\sin\theta(\bm{\pi}^{(k)},\hat{\bm{\pi}}^{(k)})\|=\mathcal{O}\!\left(\frac{C^{-1/2}}{\mathrm{gap}_{i}}\right).$ (A12)
where $\theta(\bm{\pi}^{(k)},\hat{\bm{\pi}}^{(k)})\in\mathbb{R}$ denotes the angle between the true ($\bm{\pi}^{(i)}$) and estimated ($\bm{\hat{\pi}}^{(i)}$) modes of the transfer operator. The invariant distribution is the left eigenvector at eigenvalue $1$. Perturbation theory thus implies
$\|\hat{\bm{\pi}}-\bm{\pi}\|=\mathcal{O}\!\left(\frac{C^{-1/2}}{1-|\lambda_{2}|}\right),$ (A13)
up to conditioning constants, where $1-|\lambda_{2}|$ is the mixing gap. In the small-error regime, the KL divergence is locally quadratic, so
$D_{\mathrm{KL}}(\bm{\pi}\|\hat{\bm{\pi}})=\mathcal{O}(C^{-1}).$ (A14)
Hence improved invariant-measure estimation is a direct consequence of improved finite-context operator estimation.